During startup of a power plant, a turbine accelerates from rest at 0.57 rad/s^2. How long does it take to reach its 2600 rpm operating speed? How many revolutionsdoes it make during this time?
Answer #1
final speed is 2600 rpm
1 rpm = (2π/60) rad /s
so    V =272.27 rad /s
we have speed = acceleration * time
           so time =speed/ acceleration =477.66 sec =7.96min
s=1/2 (0.57)(477.66)2
=65025.33 rad
number of revolutions = s /(2π)=10349.10 revolutions